The drive had to stop suddenly. The cars shock absorber had an elastic potential energy of 258J The spring constant of the shock absorber is 45300 N/m. Calculate the extension of the shock absorber


1
Expert's answer
2021-01-21T13:22:25-0500

Elastic Potential Energy = "\frac{1}{2}kx^2"


258"J= \frac{1}{2} \times 4530\times x^2"


"x= \sqrt{\frac{258\times 2}{4530}} = 0.3375 m"